Find the area bounded by y=2; y=√x+2; and y=2-x.


1
Expert's answer
2021-03-19T12:37:09-0400

"y=\ \sqrt x+2=>x=\left(y-2\right)^2=y^2-4y+4"

"y=2-x=>x=2-y"

"y^2-4y+4=2-y"

"y^2-3y+2=0"

"\left(y-1\right)\left(y-2\right)=0"

"y=1\ or\ y=2"

"\int_{1}^{2}{2-y-\left(y^2-4y+4\right)\ dy}"

"-2y+\frac{3}{2}y^2-\frac{y^3}{3}"

"-4+\frac{12}{2}-\frac{8}{3}-\left(-2+\frac{3}{2}-\frac{1}{3}\right)=\frac{1}{6}"